OpenStudy (anonymous):

PLZ HELP!!! Will medal ! What is the largest positive integer n$such that 1754 and 368 leave the same remainder when divided by n?

OpenStudy (is3535):

and lastly 1386=1089+297 so gcd(297,1089,1386)=gcd(297,1089)=32×11=99

OpenStudy (is3535):

99 is your answer

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

oops sorry thats not the right answer